OpenStudy (anonymous):

Simplify: (x^2-y^2)/(x+y)

mathslover (mathslover):

Use algebraic identity : \(a^2 - b^2 = (a+b)(a-b)\) So, \(x^2 - y^2 = (x+y)(x-y)\) \(\cfrac{(x^2-y^2)}{(x+y)} = \cfrac{(x+y)(x-y)}{(x+y)} \) Cancel out the like terms.